OpenStudy (anonymous):

Please help me!!! Determine the end behavior of the graph of the polynomial function: y=3x⁴+6x³-x²+12

OpenStudy (anonymous):

the degree is even (it is 4) think about the "end behaviour" of \(y=x^2\) this will be the same

OpenStudy (anonymous):

Thanks for the degree I overlooked that :p, I'm still a little confused, I'm sure its super easy. Is the end behavior seen in the equation?

OpenStudy (anonymous):



OpenStudy (abdullahm):

You might want to refresh on the end behavior concepts. http://www.purplemath.com/modules/polyends.htm The first two pictures will help you.

OpenStudy (anonymous):

as \(x\to \infty\) you have \(y\to \infty\) and as \(x\to \infty\) you have \(y\to \infty\) as wll

OpenStudy (anonymous):

Okay,thank you!

OpenStudy (abdullahm):

It was my pleasure!